# Enhancing Fetal Plane Classification Accuracy with Data Augmentation Using Diffusion Models

### v1

@@ 1 INTRODUCTION
Over the past decade, deep learning has achieved great success in medical image diagnostics, including tasks such as nodule detection Pehrson et al. (2019), image segmentation Wang et al. (2022) and classification Cai et al. (2020). Deep learning models, such as convolutional neural networks (CNNs) LeCun et al. (1989), and residual networks (ResNets) He et al. (2016), trained on large datasets, can learn intricate patterns in medical images, improving diagnostic accuracy and supporting healthcare professionals. However, the successful application of machine learning to medical imaging is heavily dependent on the availability of large, high-quality annotated datasets Alzubaidi et al. (2021). In many fields of medicine, obtaining such datasets can be difficult, and this issue is especially pronounced in ultrasound imaging Huang et al. (2021).
@@
Ultrasound imaging plays a crucial role in prenatal care, providing valuable and non-invasive insights into fetal health and development Brattain et al. (2018). Despite their widespread use, one of the key challenges in developing deep learning models for ultrasound imaging is data scarcity. There are several reasons why obtaining a large dataset of labeled ultrasound images is difficult.
@@
Firstly, due to privacy concerns, hospitals and medical institutions are often unable to release patient data, making it challenging to build large public datasets DuMont Schütte et al. (2021). Furthermore, even if hospitals are willing to share data, obtaining labeled data is costly and time-consuming. Expert knowledge is required to correctly annotate medical images, and for ultrasound, this usually means that radiologists or sonographers must manually label features of interest, such as fetal planes or abnormalities. This process is not only slow but also expensive, which limits the amount of labeled ultrasound data available for training machine learning models.
@@
Given these challenges, there has been growing interest in the use of synthetic data and generative models to supplement real-world data. Recent advances in generative models, such as Generative Adversarial Networks (GANs) Goodfellow et al. (2014) and Diffusion Models Ho et al. (2020), have opened up new possibilities for creating realistic synthetic medical images. These models can learn complex data distributions from real images and generate new synthetic images that resemble the originals. This approach has shown promise in a variety of medical imaging applications. For instance, Frid-Adar, M., et al. (2018) Frid-Adar et al. (2018) demonstrated the use of GANs for generating synthetic liver lesions in CT images. Motamed et al. (2017) Motamed et al. (2021) showed the effectiveness of using GANs to generate synthetic chest X-ray images for training classifiers. Yi et al. (2019) Yi et al. (2019) provided a comprehensive review of GAN-based techniques for generating synthetic medical images. Pan et al. (2022) Pan et al. (2023) demonstrated that diffusion models generate more realistic synthetic CT and MRI images for medical image segmentation than GANs. Hardy et al. (2023) Hardy et al. (2023) showed that generating realistic synthetic liver ultrasound images with a diffusion model can enhance liver disease classification performance. However, the application of generative models to ultrasound imaging still remains less explored, particularly for fetal ultrasound, where data scarcity is a significant challenge.
@@
In this work, we propose the use of diffusion models to generate synthetic fetal ultrasound images and investigate their potential to improve the performance of machine learning classifiers. Unlike GANs, which can suffer from training instability and mode collapse Kodali et al. (2017), diffusion models offer more stable training dynamics and produce high-quality, diverse samples, making them well-suited for this application Dhariwal and Nichol (2021). Specifically, we hypothesize that combining real ultrasound images with generated images can enhance classifier accuracy, especially in cases where annotated real data is limited. By training the model first on synthetic images and then fine-tuning it on real images, we aim to overcome the data scarcity problem in fetal ultrasound imaging. We make the following key contributions:
@@ 2.1 Traditional Approaches
Traditionally, data scarcity in medical imaging has been addressed through various data augmentation techniques. These methods aim to artificially expand the size of a dataset without requiring additional labeled images. Early data augmentation strategies for medical images included transformations such as rotation, scaling, cropping, flipping, and elastic deformations. These techniques were particularly useful in fields like MRI and CT imaging, where the generation of new images from the same subject could help the model generalize better Goceri (2023). For general imaging, methods such as image cropping and noise addition have been widely used. Simple geometric transformations, such as rotation and scaling have been applied broadly as well. While these methods can be effective in increasing the volume of data, they often fail to generate truly novel variations of the data, which limits their ability to capture the full diversity of possible real-world cases Goceri (2023).
Another traditional approach to tackling data scarcity is transfer learning, which has been used in many medical imaging tasks Sanford et al. (2020). Transfer learning allows models trained on large, publicly available datasets (such as ImageNet) to be adapted to medical datasets with fewer labeled samples. While effective in some contexts, transfer learning still relies heavily on the availability of a reasonable amount of high-quality annotated data, making it less suitable in situations where data is extremely limited or highly specialized, such as in fetal ultrasound imaging.
@@ 2.2 Deep Learning-Based Approaches
In particular, Generative Adversarial Networks (GANs) have been explored as a way to generate synthetic medical images. GANs consist of two neural networks: a generator that creates synthetic data, and a discriminator that attempts to distinguish between real and fake data. GANs have been applied to various imaging modalities, including CT, MRI, and ultrasound. For instance, Frid-Adar et al. (2018) Frid-Adar et al. (2018) used GANs to generate synthetic CT images of liver lesions, which were used to augment a classifier for liver lesion detection. Similarly, Motamed et al. (2017) Motamed et al. (2021) showed the effectiveness of using GANs to generate synthetic chest X-ray images for training classifiers. In ultrasound imaging, GANs have shown promise in generating synthetic images for various purposes. Montero et al. (2021) Montero et al. (2021) applied a GAN-based approach to generate synthetic fetal ultrasound images, focusing on augmenting data for training classifiers on fetal development stages. While GANs have shown considerable success, one challenge with these models is the difficulty in training them, especially in medical imaging, where the distribution of data can be complex and diverse. In addition to GANs, Variational Autoencoders (VAEs) have also been used for generating synthetic medical images. VAEs are another class of generative models that learn a probabilistic mapping between a low-dimensional latent space and high-dimensional image space. Studies such as Dorent et al. (2023) Dorent et al. (2023) have demonstrated the potential of VAEs for generating synthetic ultrasound and MRI images for the purposes of data augmentation.
@@
In recent years, diffusion models have emerged as a powerful new class of generative models, particularly in the field of image synthesis Ho et al. (2020). Diffusion models work by simulating a gradual “noising” process that adds random noise to data in a series of steps, and then by learning how to reverse this process they can generate high-quality samples. One key advantage of diffusion models over GANs and VAEs is their ability to produce high-fidelity images with more stable training dynamics.
@@
The use of diffusion models in medical imaging is a relatively recent development, but promising results have already been reported in various applications. Dhariwal and Nichol (2021) Dhariwal and Nichol (2021) demonstrated the use of diffusion models for generating high-quality images in a variety of domains, including natural images, and showed that they outperform GANs in terms of image quality and diversity. In medical imaging, recent studies have started to explore diffusion models for generating synthetic data. Pan et al. (2022) Pan et al. (2023) applied diffusion models to generate synthetic CT and MRI images for medical image segmentation tasks, showing that diffusion models can generate more realistic images compared to traditional methods like GANs. Similarly, in the field of ultrasound imaging, Hardy et al. (2023) Hardy et al. (2023) used a diffusion model to generate realistic synthetic liver ultrasound images, demonstrating that such generated data could improve the performance of a liver disease classification model. However, the application of diffusion models to fetal ultrasound images remains an underexplored area, and there is a gap in the literature regarding how these models can be applied to enhance classification performance for fetal health assessment.
@@ 3.1 Classifier-Guided Diffusion Model
We use a classifier-guided diffusion model to generate synthetic ultrasound images Dhariwal and Nichol (2021). In this approach, the diffusion model is conditioned on the classifier’s predictions, allowing it to generate images that belong to specific classes. This method enhances the model’s ability to produce class-specific images that better align with real-world data, improving the quality and relevance of the generated samples. The overall process consists of two interconnected models: the diffusion model and the classifier.
@@ 3.2 Ultrasound Image Dataset
To train and evaluate our proposed method, we use the FETAL_PLANES_DB Burgos-Artizzu et al. (2020), a commonly used maternal-fetal ultrasound image dataset. The FETAL_PLANES_DB contains in total $12400$ images of fetal planes, which are essential in prenatal care for monitoring fetal development. The dataset is divided into six distinct classes, i.e., Abdomen, Brain, Femur, Thorax, the mother’s cervix (widely used for prematurity screening), and a general category to include any other less common image plane. These classes represent different features that are crucial for assessing fetal health.
@@ 4 EXPERIMENTS
In this section, we describe the details of the experimental setup used to evaluate the proposed method. This includes the dataset acquisition and preparation process, data preprocessing, the architecture of the models used, and the techniques employed for image generation and classification. We also provide details on the evaluation metrics used to assess the quality of the generated images and the classification performance.
@@ 4.1 Dataset Acquisition and Preparation
The FETAL_PLANES_DB is a publicly available maternal-fetal ultrasound image dataset https://zenodo.org/records/3904280. As shown in Table 1, the FETAL_PLANES_DB dataset was split into an $80/20$ training and test set, resulting in $9923$ images for training and $2477$ images for testing. This common split ensures a representative sample of the data is used to train the model while reserving a separate set of images to evaluate its generalization ability on unseen data. The $80\%$ portion (training set) is used to train the model. The model learns patterns and relationships within the data to make accurate predictions on new data. The $20\%$ portion (test set) was used to assess the model’s performance on unseen data. This helped identify how well the model generalized to real-world scenarios beyond the training data. By splitting the data in this way, a balance between training the model effectively could be achieved and ensuring it performed well on data it has not encountered before.
@@ 4.2 Data Preprocessing
Before training the models, several preprocessing steps on the real images were performed. First, all images were normalized to have pixel values in the range $[0,1]$. This ensures that the neural networks can learn efficiently by reducing the risk of vanishing or exploding gradients. Second, the ultrasound images in the dataset were resized to a uniform resolution of $128\times 128$ pixels. This size was chosen to maintain sufficient detail in the images while reducing computational overhead. Each image in the dataset was assigned one of the six class labels corresponding to different fetal anatomical planes. The synthetic images generated by the diffusion model were also labelled according to their respective class. We also applied traditional data augmentation techniques to real images. Random rotations of the images between $-30^{\circ}$ and $30^{\circ}$ was employed to simulate different scanning angles. Horizontal and vertical flippings were used to account for different fetal orientations in the ultrasound.

### v2

@@ 1 INTRODUCTION
Over the past decade, deep learning has achieved great success in medical image diagnostics, including tasks such as nodule detection (Pehrson et al., 2019), image segmentation (Wang et al., 2022) and classification (Cai et al., 2020). Deep learning models, such as convolutional neural networks (CNNs) (LeCun et al., 1989), and residual networks (ResNets) (He et al., 2016), trained on large datasets, can learn intricate patterns in medical images, improving diagnostic accuracy and supporting healthcare professionals. However, the successful application of machine learning to medical imaging is heavily dependent on the availability of large, high-quality annotated datasets (Alzubaidi et al., 2021). In many fields of medicine, obtaining such datasets can be difficult, and this issue is especially pronounced in ultrasound imaging (Huang et al., 2021).
@@
Ultrasound imaging plays a crucial role in prenatal care, providing valuable and non-invasive insights into fetal health and development (Brattain et al., 2018). Despite their widespread use, one of the key challenges in developing deep learning models for ultrasound imaging is data scarcity. There are several reasons why obtaining a large dataset of labeled ultrasound images is difficult.
@@
Firstly, due to privacy concerns, hospitals and medical institutions are often unable to release patient data, making it challenging to build large public datasets (DuMont Schütte et al., 2021). Furthermore, even if hospitals are willing to share data, obtaining labeled data is costly and time-consuming. Expert knowledge is required to correctly annotate medical images, and for ultrasound, this usually means that radiologists or sonographers must manually label features of interest, such as fetal planes or abnormalities. This process is not only slow but also expensive, which limits the amount of labeled ultrasound data available for training machine learning models.
@@
Given these challenges, there has been growing interest in the use of synthetic data and generative models to supplement real-world data. Recent advances in generative models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Diffusion Models (Ho et al., 2020), have opened up new possibilities for creating realistic synthetic medical images. These models can learn complex data distributions from real images and generate new synthetic images that resemble the originals. This approach has shown promise in a variety of medical imaging applications. For instance, Frid-Adar et al. (2018) demonstrated the use of GANs for generating synthetic liver lesions in CT images. Motamed et al. (2021) showed the effectiveness of using GANs to generate synthetic chest X-ray images for training classifiers. Yi et al. (2019) provided a comprehensive review of GAN-based techniques for generating synthetic medical images. Pan et al. (2023) demonstrated that diffusion models generate more realistic synthetic CT and MRI images for medical image segmentation than GANs. Hardy et al. (2023) showed that generating realistic synthetic liver ultrasound images with a diffusion model can enhance liver disease classification performance. More recently, Qiu et al. (2025) proposed a noise-consistent Siamese-Diffusion model for medical image synthesis and segmentation, further expanding the capabilities of diffusion models in this domain. However, the application of generative models to ultrasound imaging still remains less explored, particularly for fetal ultrasound, where data scarcity is a significant challenge.
@@
In this work, we propose the use of diffusion models to generate synthetic fetal ultrasound images and investigate their potential to improve the performance of machine learning classifiers. Unlike GANs, which can suffer from training instability and mode collapse (Kodali et al., 2017), diffusion models offer more stable training dynamics and produce high-quality, diverse samples, making them well-suited for this application (Dhariwal and Nichol, 2021). Specifically, we hypothesize that combining real ultrasound images with generated images can enhance classifier accuracy, especially in cases where annotated real data is limited. By training the model first on synthetic images and then fine-tuning it on real images, we aim to overcome the data scarcity problem in fetal ultrasound imaging. We make the following key contributions:
@@ 2.1 Traditional Approaches
Traditionally, data scarcity in medical imaging has been addressed through various data augmentation techniques. These methods aim to artificially expand the size of a dataset without requiring additional labeled images. Early data augmentation strategies for medical images included transformations such as rotation, scaling, cropping, flipping, and elastic deformations (Hussain et al., 2017). These techniques were particularly useful in fields like MRI and CT imaging, where the generation of new images from the same subject could help the model generalize better (Goceri, 2023). For general imaging, methods such as image cropping and noise addition have been widely used. Simple geometric transformations, such as rotation and scaling have been applied broadly as well. While these methods can be effective in increasing the volume of data, they often fail to generate truly novel variations of the data, which limits their ability to capture the full diversity of possible real-world cases (Goceri, 2023).
Another traditional approach to tackling data scarcity is transfer learning, which has been used in many medical imaging tasks (Sanford et al., 2020). Transfer learning allows models trained on large, publicly available datasets (such as ImageNet) to be adapted to medical datasets with fewer labeled samples. While effective in some contexts, transfer learning still relies heavily on the availability of a reasonable amount of high-quality annotated data, making it less suitable in situations where data is extremely limited or highly specialized, such as in fetal ultrasound imaging.
@@ 2.2 Deep Learning-Based Approaches
In particular, Generative Adversarial Networks (GANs) have been explored as a way to generate synthetic medical images. GANs consist of two neural networks: a generator that creates synthetic data, and a discriminator that attempts to distinguish between real and fake data. GANs have been applied to various imaging modalities, including CT, MRI, and ultrasound. For instance, Frid-Adar et al. (2018) used GANs to generate synthetic CT images of liver lesions, which were used to augment a classifier for liver lesion detection. Similarly, Motamed et al. (2021) showed the effectiveness of using GANs to generate synthetic chest X-ray images for training classifiers. In ultrasound imaging, GANs have shown promise in generating synthetic images for various purposes. Montero et al. (2021) applied a GAN-based approach to generate synthetic fetal ultrasound images, focusing on augmenting data for training classifiers on fetal development stages. While GANs have shown considerable success, one challenge with these models is the difficulty in training them, especially in medical imaging, where the distribution of data can be complex and diverse. In addition to GANs, Variational Autoencoders (VAEs) have also been used for generating synthetic medical images. VAEs are another class of generative models that learn a probabilistic mapping between a low-dimensional latent space and high-dimensional image space. Studies such as Dorent et al. (2023) have demonstrated the potential of VAEs for generating synthetic ultrasound and MRI images for the purposes of data augmentation.
@@
In recent years, diffusion models have emerged as a powerful new class of generative models, particularly in the field of image synthesis (Ho et al., 2020). Diffusion models work by simulating a gradual “noising” process that adds random noise to data in a series of steps, and then by learning how to reverse this process they can generate high-quality samples. One key advantage of diffusion models over GANs and VAEs is their ability to produce high-fidelity images with more stable training dynamics.
@@
The use of diffusion models in medical imaging is a relatively recent development, but promising results have already been reported in various applications. Dhariwal and Nichol (2021) demonstrated the use of diffusion models for generating high-quality images in a variety of domains, including natural images, and showed that they outperform GANs in terms of image quality and diversity. In medical imaging, recent studies have started to explore diffusion models for generating synthetic data. Pan et al. (2023) applied diffusion models to generate synthetic CT and MRI images for medical image segmentation tasks, showing that diffusion models can generate more realistic images compared to traditional methods like GANs. Similarly, in the field of ultrasound imaging, Hardy et al. (2023) used a diffusion model to generate realistic synthetic liver ultrasound images, demonstrating that such generated data could improve the performance of a liver disease classification model. However, the application of diffusion models to fetal ultrasound images remains an underexplored area, and there is a gap in the literature regarding how these models can be applied to enhance classification performance for fetal health assessment.
2.3 Fetal Plane Classification
Recent advancements in fetal ultrasound classification have explored diverse approaches. For instance, Krishna and Kokil (2023) utilized a multi-layer perceptron with deep feature integration to automate the classification of maternal fetal ultrasound planes. This work demonstrated the potential of integrating deep features for improved classification accuracy. Similarly, Krishna and Kokil (2024a) employed a stacked ensemble of deep learning models, achieving high performance through the combination of multiple classifiers. However, our method focuses on the performance of a single network enhanced by diffusion models. Also, Rauf et al. (2023) proposed a deep bottleneck residual 82-layered architecture with Bayesian optimization for ultrasound plane classification. This method differs from our work by its focus on architectural optimization and Bayesian methods, whereas we focus on data augmentation through diffusion models. Transfer learning has also been explored, as seen in Ghabri et al. (2023), which focused on accurate fetal organ classification, showing the potential of pre-trained models in this domain. Furthermore, the works of Krishna and Kokil (2024b, c) specifically address the automated identification of fetal biometry planes, utilizing techniques such as adaptive channel weighting. These works highlight the importance of accurate biometry plane identification, which is a key step in fetal assessment. However, they do not explore the use of diffusion models for data augmentation, as we do in our work.
@@ 3 METHODS
3.1 Basics of Diffusion Models
Diffusion models are a class of generative models that have recently shown remarkable success in synthesizing high-quality images. Unlike Generative Adversarial Networks (GANs), which learn to generate images through an adversarial process, diffusion models operate by progressively adding noise to data and then learning to reverse this noisy process to generate new data. The core idea is inspired by non-equilibrium thermodynamics.
The Forward (Diffusion) Process. In the forward process, a data sample $x_{0}$ (e.g., an image) is gradually corrupted by adding Gaussian noise over a series of $T$ timesteps. At each timestep $t$, a small amount of noise is added to the previous timestep’s sample $x_{t-1}$ to produce $x_{t}$. This process is defined such that the distribution $q(x_{t}|x_{t-1})$ is a Gaussian, and consequently, the distribution $q(x_{t}|x_{0})$ can be directly sampled:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$
where $\alpha_{t}$ is a predefined variance schedule, typically decreasing over time, such that $x_{T}$ eventually becomes pure Gaussian noise.
The Reverse (Denoising) Process. The goal of a diffusion model is to learn the reverse process, which involves denoising a noisy sample $x_{t}$ back to a cleaner $x_{t-1}$, and eventually to the original data $x_{0}$. This reverse process is also modeled as a Gaussian, but with learned means and variances:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are learned parameters of a neural network.
Training Objective. Dhariwal and Nichol (2021) builds upon the denoising diffusion probabilistic models (DDPMs) framework, where the training objective for diffusion models is to maximize the likelihood of the training data. This is achieved by optimizing a variational lower bound (VLB) on the negative log-likelihood.
A simplified and commonly used training objective, particularly in the context of the mentioned paper, is to train a neural network (often denoted as $\epsilon_{\theta}$) to predict the noise that was added at each timestep $t$. Specifically, the objective function is a mean squared error between the predicted noise and the actual noise.
For a given noisy sample $x_{t}$ at timestep $t$, obtained by adding noise $\epsilon\sim\mathcal{N}(0,I)$ to $x_{0}$:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon$
The neural network $\epsilon_{\theta}(x_{t},t)$ is trained to predict $\epsilon$. The simplified training objective can be expressed as:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t\sim[1,T],x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\theta}(x_{t},t)\right\|^{2}\right]$
where:
  • $t$ is sampled uniformly from $1$ to $T$.
  • $x_{0}$ is a data sample from the true data distribution $q(x_{0})$.
  • $\epsilon$ is the pure Gaussian noise sampled from $\mathcal{N}(0,I)$.
  • $x_{t}$ is the noisy sample at timestep $t$, obtained using the forward process formula above.
  • $\epsilon_{\theta}(x_{t},t)$ is the neural network’s prediction of the noise, conditioned on $x_{t}$ and the timestep $t$.
  • $||\cdot||^{2}$ denotes the squared L2 norm.
This objective aims to minimize the difference between the true noise $\epsilon$ and the noise predicted by the model $\epsilon_{\theta}$. By minimizing this objective, the model learns to accurately predict the noise component at each timestep, which in turn allows it to effectively reverse the diffusion process and generate high-quality samples.
@@ 3.2 Classifier-Guided Diffusion Model
We use a classifier-guided diffusion model to generate synthetic ultrasound images (Dhariwal and Nichol, 2021). In this approach, the diffusion model is conditioned on the classifier’s predictions, allowing it to generate images that belong to specific classes. This method enhances the model’s ability to produce class-specific images that better align with real-world data, improving the quality and relevance of the generated samples. The overall process consists of two interconnected models: the diffusion model and the classifier.
@@ 3.3 Ultrasound Image Dataset
To train and evaluate our proposed method, we use the FETAL_PLANES_DB (Burgos-Artizzu et al., 2020), a commonly used maternal-fetal ultrasound image dataset. The FETAL_PLANES_DB contains in total $12400$ images of fetal planes, which are essential in prenatal care for monitoring fetal development. The dataset is divided into six distinct classes, i.e., Abdomen, Brain, Femur, Thorax, the mother’s cervix (widely used for prematurity screening), and a general category to include any other less common image plane. These classes represent different features that are crucial for assessing fetal health.
@@ 4 EXPERIMENTS
In this section, we describe the details of the experimental setup used to evaluate the proposed method. This includes the dataset acquisition and preparation process, data preprocessing, the architecture of the models used, and the techniques employed for image generation and classification. We also provide details on the evaluation metrics used to assess the quality of the generated images and the classification performance. All experiments in the original submission were conducted using a single Quadro RTX 5000 GPU with 16GB memory. Most experiments in the revised version were conducted using 8 A100 GPUs with 40GB memory.
@@ 4.1 Dataset Acquisition and Preparation
The FETAL_PLANES_DB is a publicly available maternal-fetal ultrasound image dataset https://zenodo.org/records/3904280. As shown in Table 1, the FETAL_PLANES_DB dataset was randomly split into an $80/20$ training and test set, resulting in $9923$ images for training and $2477$ images for testing. The detailed training and testing dataset splitting ratio for each class can also be found in Table 1. This common split ensures a representative sample of the data is used to train the model while reserving a separate set of images to evaluate its generalization ability on unseen data. The $80\%$ portion (training set) is used to train the model. The model learns patterns and relationships within the data to make accurate predictions on new data. The $20\%$ portion (test set) was used to assess the model’s performance on unseen data. This helped identify how well the model generalized to real-world scenarios beyond the training data. By splitting the data in this way, a balance between training the model effectively could be achieved and ensuring it performed well on data it has not encountered before.
@@ 4.2 Data Preprocessing
Before training the models, several preprocessing steps on the real images were performed. First, all images were normalized to have pixel values in the range $[0,1]$. This ensures that the neural networks can learn efficiently by reducing the risk of vanishing or exploding gradients. Second, the ultrasound images in the dataset were resized to a uniform resolution of $128\times 128$ pixels. This size was chosen to maintain sufficient detail in the images while reducing computational overhead. Each image in the dataset was assigned one of the six class labels corresponding to different fetal anatomical planes. The synthetic images generated by the diffusion model were also labelled according to their respective class. We also applied traditional data augmentation techniques to real images. Random rotations of the images between $-30^{\circ}$ and $30^{\circ}$ were employed to simulate different scanning angles. Horizontal and vertical flippings were used to account for different fetal orientations in the ultrasound.